What is the percentage by mass of solution formed by dissolving 27.0 grams of HCl in 102.0 grams of water?


1
Expert's answer
2021-05-28T00:56:00-0400

Total mass = 27.0+102.0 = 129.0 g

Proportion:

129.0 – 100%

27.0 – x

"x = \frac{27.0 \times 100 \; \%}{129.0} = 20.93 \; \%"

Answer: 20.93 %